### v1

# SSL4RL: Revisiting Self-supervised Learning as Intrinsic Reward for Visual-Language Reasoning

Abstract
Vision-language models (VLMs) have shown remarkable abilities by integrating large language models with visual inputs. However, they often fail to utilize visual evidence adequately, either depending on linguistic priors in vision-centric tasks or resorting to textual shortcuts during reasoning. Although reinforcement learning (RL) can align models with desired behaviors, its application to VLMs has been hindered by the lack of scalable and reliable reward mechanisms. To overcome this challenge, we propose SSL4RL, a novel framework that leverages self-supervised learning (SSL) tasks as a source of verifiable rewards for RL-based fine-tuning. Our approach reformulates SSL objectives—such as predicting image rotation or reconstructing masked patches—into dense, automatic reward signals, eliminating the need for human preference data or unreliable AI evaluators. Experiments show that SSL4RL substantially improves performance on both vision-centric and vision-language reasoning benchmarks. Furthermore, through systematic ablations, we identify key factors—such as task difficulty, model scale, and semantic alignment with the target domain—that influence the effectiveness of SSL4RL tasks, offering new design principles for future work. We also demonstrate the framework’s generality by applying it to graph learning, where it yields significant gains. SSL4RL establishes a versatile and effective paradigm for aligning multimodal models using verifiable, self-supervised objectives.
1 Introduction
Vision–Language Models (VLMs) have rapidly advanced multimodal understanding by leveraging the expert-level reasoning capabilities of Large Language Models (LLMs). This synergy has enabled broad applications, from visual question answering to interactive dialogue. Yet the reliance on linguistic priors introduces systematic weaknesses. For vision-centric tasks—where answers must be derived solely from image content, such as classification—VLMs often lag behind specialist vision models (Fu et al., 2025; Tong et al., 2024; Fu et al., 2024). Conversely, in vision–language reasoning tasks, VLMs tend to exploit textual knowledge rather than grounding their reasoning in visual evidence, a tendency amplified in long-form generation (Jian et al., 2025). These limitations underscore the need for training methods that reinforce visual grounding and robust reasoning simultaneously.
Reinforcement learning (RL) has emerged as the dominant paradigm for post-training large models, demonstrating that preference-based signals—collected from humans or distilled from AI feedback—can substantially improve helpfulness and alignment (Ouyang et al., 2022; Rafailov et al., 2023; Bai et al., 2022). More recently, verifier-driven RL has shown particular promise: models trained with automatically checkable rewards achieve striking gains in domains such as mathematics and programming (Le et al., 2022; Shao et al., 2024; Yue et al., 2025). However, these successes expose a fundamental bottleneck: outside domains with explicit programmatic verifiers, scalable and reliable rewards are scarce. In such cases, training pipelines often revert to LLM-as-a-judge heuristics, which are biased, noisy, and prone to adversarial manipulation (Raina et al., 2024; Chen et al., 2024b). This raises a key question: how can we obtain abundant, verifiable reinforcement signals for VLMs in domains where external verifiers are absent?
Self-supervised learning (SSL) offers a natural but underexplored answer. SSL has been central to representation learning across modalities: masked language modeling and next-token prediction for text (Devlin et al., 2019; Lewis et al., 2020; Raffel et al., 2020), contrastive learning and masked autoencoders for vision (Gidaris et al., 2018; Noroozi & Favaro, 2016; Doersch et al., 2015; Chen et al., 2020; He et al., 2020; Grill et al., 2020; Caron et al., 2021; He et al., 2022). The principle is simple yet powerful: perturb data, and require the model to reconstruct or discriminate. Crucially, SSL tasks define intrinsically verifiable targets. Given an image and its rotated variant, the ground-truth angle is unambiguous. Unlike preference-based signals, these targets are properties of the data itself, providing dense, reliable supervision at scale.
Figure 1: Overview of the SSL4RL framework. A corruption function transforms an input into a context–target pair. The model conditions on the context, generates predictions, and receives a verifiable reward by comparing against the target. The reward is then used to optimize the model via Reinforcement Learning (RL).
In this paper, we propose SSL4RL, a general framework that repurposes SSL tasks as verifiable reward functions for RL-based post-training. Instead of treating SSL solely as a pre-training tool (Tong et al., 2024), we reinterpret it through the lens of RL: corrupted inputs define trajectories, correctness defines rewards, and policy optimization drives updates. SSL4RL requires no human labels, external verifiers, or heuristic judges, yet produces dense and scalable reinforcement signals. Importantly, unlike conventional SSL, SSL4RL emphasizes generating natural language reasoning paths to solve vision tasks, thereby bridging perceptual learning and reasoning alignment.
We systematically evaluate SSL4RL on both vision-centric and multimodal reasoning benchmarks. On ImageNet-1K, SSL4RL significantly improves classification accuracy over the base model. For vision–language reasoning tasks, it delivers consistent gains, with average improvements of 9% on MMBench and 8% on SEED-Bench. A key finding is that the effectiveness of SSL tasks in SSL4RL differs from their role in traditional SSL: contrastive objectives, while dominant in pre-training, show limited benefit unless paired with stronger data augmentation, whereas position prediction—often deemed too trivial for SSL—emerges as highly effective in the SSL4RL setting. The ablation study reveals that the efficacy of an SSL4RL task is strongly influenced by model capacity, its semantic alignment with downstream tasks, and the inherent difficulty of the task itself. These findings offer initial insights into what makes a good SSL4RL task.
We demonstrate that the SSL4RL paradigm generalizes beyond vision by extending it to the graph domain. Through three graph-related SSL tasks—attribute masking, neighbor prediction, and link prediction (Hou et al., 2022; Hu et al., 2020)—we achieve marked improvements on node classification and link prediction benchmarks. These findings highlight SSL4RL as a versatile recipe for extracting verifiable rewards with self-supervised tasks.
Contributions. Our work makes two key contributions: (1) we introduce SSL4RL, a unified framework bridging self-supervised learning and RL-based post-training through verifiable rewards; (2) we provide a comprehensive cross-domain study identifying which SSL tasks best transfer to reasoning and which do not. Together, these results challenge the assumption that all self-supervision is equally useful, and emphasize that the abundance of verifiable signals in SSL can be harnessed not only for representation learning, but also to drive alignment and reasoning in VLMs.
2 Related Work
RL training with external verifiers. Reinforcement learning has become a dominant paradigm for aligning large models. RLHF aligns LLMs with human intent through preference data (Ouyang et al., 2022), while Direct Preference Optimization (DPO) reframes preference learning as a contrastive loss without explicit reward models (Rafailov et al., 2023). Constitutional AI replaces human raters with rule-based AI feedback (Bai et al., 2022). More recently, verifier-driven RL has achieved notable success in domains such as code and math, where correctness can be mechanically checked (Le et al., 2022; Shao et al., 2024). However, in domains lacking such verifiers, many systems fall back on LLM-as-a-judge signals. While convenient, these rewards are biased, noisy, and adversarially manipulable (Raina et al., 2024; Chen et al., 2024b), undermining their reliability. This limitation motivates the search for scalable alternatives where correctness is intrinsic to the data.
RL training with self-reward. Several methods attempt to reduce dependence on external labels or verifiers by enabling models to generate their own training signals. Self-Instruct (Wang et al., 2023b) bootstraps fine-tuning data through synthetic instructions, while STaR (Zelikman et al., 2022) and Reflexion (Shinn et al., 2023) improve reasoning via self-generated rationales and feedback. Self-Consistency (Wang et al., 2023a) aggregates multiple sampled rationales to increase reliability. Reinforced Pre-Training (RPT) scales this idea by using RL objectives such as next-token prediction at the pretraining stage (Liu et al., 2025). While these methods reduce the need for human labels, they still optimize toward approximating original task accuracy, often requiring correctness evaluation or model-judging heuristics. In contrast, our approach seeks verifiable, abundant signals outside the original task by reinterpreting SSL tasks as reinforcement rewards.
Self-supervised learning across modalities. Self-supervision has been the foundation of representation learning across domains. In language, pretext tasks include masked language modeling (Devlin et al., 2019; Lewis et al., 2020; Raffel et al., 2020) and next-token prediction. Variants that mask reasoning steps show promise in enhancing mathematical reasoning (Chen et al., 2024a). In vision, early tasks include rotation (Gidaris et al., 2018), jigsaw (Noroozi & Favaro, 2016), and context prediction (Doersch et al., 2015), while modern SSL emphasizes contrastive learning (Chen et al., 2020; He et al., 2020; Grill et al., 2020; Caron et al., 2021) and generative masking (He et al., 2022). In multimodal learning, CLIP (Radford et al., 2021) and ALBEF (Li et al., 2021) demonstrate the power of contrastive and distillation objectives for aligning vision and language. Graph SSL extends these principles to structural data, with node/edge masking (Hu et al., 2020; Hou et al., 2022) and contrastive augmentations (You et al., 2020). A unifying property of all these objectives is their intrinsically verifiable targets, making them natural candidates for repurposing as RL rewards.
Building on these threads, our work unifies SSL and RL post-training. By treating SSL objectives as reinforcement rewards, SSL4RL supplies dense, scalable, and verifiable signals without human annotations, model-judging heuristics, or handcrafted verifiers. This perspective highlights that the supervision already embedded in SSL tasks can be harnessed to drive reasoning improvements.
3 SSL4RL Framework
We formalize SSL4RL as a general recipe for converting self-supervised learning (SSL) objectives into reinforcement learning (RL) rewards for post-training large language models and related architectures. This section introduces the notation, shows how SSL tasks are reinterpreted under the RL formalism, and describes the optimization strategy. A high-level illustration of the framework is shown in Figure 1.
Problem Setup. Let $\pi_{\theta}$ denote a parametric model with parameters $\theta$, defined over sequences of actions. In vision–language tasks, actions may include discrete classification labels (e.g., rotations, patch indices) or text tokens. We assume access to a data distribution $\mathcal{D}$ of inputs $x\in\mathcal{X}$, which can be text and images. In the standard RL formalism, a trajectory $\tau$ is generated by rolling out $\pi_{\theta}$ in an environment $\mathcal{E}$, and receives a scalar reward $R(\tau)$. In SSL4RL, the “environment” is defined by a corruption function $c(x)=(\tilde{x},y)$, which maps an input $x$ into a corrupted context $\tilde{x}$ and a ground-truth target $y$. The policy $\pi_{\theta}$ conditions on $\tilde{x}$ to produce an output $\hat{y}$, and a reward $r(\hat{y},y)$ is computed based on agreement with the ground truth. Thus, every SSL task induces an RL task: the corruption defines the environment, the target defines the verifiable ground truth, and the reward function provides supervision.
Figure 2: Four SSL4RL tasks considered in our study. Rotation: An image is rotated by a predefined angle, and the task is to predict this angle. Jigsaw: After dividing an image into a grid and permuting the patches, the goal is to predict the correct permutation index. Contrastive: Two augmented views are generated from an image, and the objective is to identify whether two given views originate from the same source image. Position: Given an image and a single patch cropped from it, the task is to predict the patch’s original spatial position.
From SSL to RL Rewards. Our work revisits four representative SSL pretext tasks, depicted in Figure 2. Each SSL task is defined by a tuple $(c,y,r)$, where $c$ is a corruption function applied to the inputs $x$ (specifically referring to images), $y$ is the self-supervised target derived from the corruption, and $r$ is the reward signal computed based on the model’s prediction $\hat{y}$. Specifically,
  • Rotation Prediction (Gidaris et al., 2018): $c(x,y)$ rotates $x$ by an angle $y$. The reward is $r=\mathbb{1}[\hat{y}=y]$, where $\hat{y}$ is the predicted angle.
  • Jigsaw Puzzles (Noroozi & Favaro, 2016): $c(x,y)$ partitions $x$ into a grid and permutes patches by index $y$. The reward is $r=\mathbb{1}[\hat{y}=y]$, where $\hat{y}$ is the predicted permutation.
  • Contrastive Learning (Chen et al., 2020; Radford et al., 2021): $c(x)$ generates augmented views. The reward $r$ is a binary classification, imitating the InfoNCE similarity score that encourages high similarity for positive pairs and low similarity for negatives.
  • Patch Position Prediction (Doersch et al., 2015): $c(x,y)$ extracts a patch from location $y$. The reward is $r=\mathbb{1}[\hat{y}=y]$, where $\hat{y}$ is the predicted location.
These rewards are verifiable as they are computed against unambiguous ground-truth targets $y$.
Optimization via GRPO. We adopt Grouped Reinforcement Policy Optimization (GRPO) (Shao et al., 2024), an efficient policy-gradient method designed for large-scale LLM training. Given a reference policy $\pi_{0}$ (e.g., the supervised fine-tuned model before RL), GRPO optimizes $\pi_{\theta}$ by maximizing the following regularized objective:
$\mathcal{J}(\theta)=\mathbb{E}_{\tau\sim\pi_{\theta}}\!\left[R(\tau)-\beta\,\mathrm{KL}\!\left(\pi_{\theta}(\cdot|\tau)\,\|\,\pi_{0}(\cdot|\tau)\right)\right],$ (1)
where $R(\tau)$ is the SSL-derived reward, and $\beta$ controls the strength of KL regularization. The KL penalty prevents divergence from the reference distribution and stabilizes training. GRPO performs updates by sampling rollouts, normalizing rewards across groups, and applying clipped policy-gradient updates similar to PPO (Schulman et al., 2017), but in a manner more compute-efficient for large batch LLM training. All experiments in this work apply the same GRPO configuration across tasks, ensuring that differences in outcomes are attributable to the choice of SSL reward.
4 What Makes a Good SSL4RL Task?
In this section, we examine the core principles for designing effective SSL4RL tasks through comprehensive experiments on vision-language reasoning (Section 4.1) and vision-centric benchmarks (Section 4.3). Our results indicate that standard SSL strategies do not directly transfer to the SSL4RL setting. In Section 4.2, we present ablation studies on task difficulty scaling, model size, and task combinations to elucidate these design principles further.
4.1 Experiments on Vision-Language Reasoning Tasks
SSL Task Settings. For each SSL pretext task, we implement the following configurations. In Rotation, images are rotated counterclockwise by a randomly selected angle from 0°, 90°, 180°, 270°. In Jigsaw, each image is partitioned into a 2×2 grid, and the patches are randomly shuffled. In Contrastive, we apply the standard augmentation pipeline from Chen et al. (2020), including color jittering, grayscale conversion, Gaussian blur, horizontal flipping, and random resized cropping, each with an application probability of 0.2. In Position, the image is divided into four equal quadrants, and the target is to identify which quadrant (upper-left, upper-right, lower-left, lower-right) contains a specified patch.
Training and Evaluation Settings. Our experiments primarily adopt Qwen-2.5-VL-3B/7B-Instruct (Bai et al., 2025) for their moderate sizes and strong reasoning performance. For GRPO training, we set the group size to be 5, the KL loss coefficient to be 0.01, the entropy loss coefficient to be 0, and context length to be 2048. We train the models on 8xA800 GPUs with batch size of 512. For evaluation, we adopt the third-party evaluation tool VLMEvalKit (Duan et al., 2024), with the default sampling configured with a temperature of 0.01, top-p of 0.001, and top-k of 1.
Benchmarks. We assess our approach on two prominent multi-modal benchmarks designed for evaluating vision-language reasoning capabilities: MMBench (Liu et al., 2024) and SEED-Bench (Li et al., 2023). MMBench provides a diverse testbed with over 3,000 multiple-choice questions spanning 20 distinct ability dimensions. All results on MMBench are reported for the DEV-EN split, showing the average performance per category (detailed per-dimension results are in Appendix A). SEED-Bench is a comprehensive benchmark with human annotations for image and video modalities. For evaluation, we select the 9 core dimensions pertaining to image understanding, which comprise 14,232 examples in total.
Results. The evaluation results on MMBench and SEED-Bench are presented in Table 1 and Table 2, respectively. Overall, the proposed SSL4RL paradigm leads to substantial performance gains on both MMBench and SEED-Bench benchmarks. On average, SSL4RL-tuned models outperform the base model by 7.39% on MMBench and 8.94% on SEED-Bench. Notably, SSL4RL achieves a remarkable improvement of 39.00 percentage points (80.54% vs. 41.54%) on the Relation Reasoning task in MMBench. On the Visual Reasoning task of SEED-Bench, the improvement reaches up to 19.63 percentage points (73.41% vs. 53.78%). These consistent improvements validate the effectiveness of SSL4RL in enhancing vision-language reasoning.
Table 1: Test performance (%) on MMBench downstream tasks. Logical: Logical Reasoning, Relation: Relation Reasoning, Attribute: Attribute Reasoning, Coarse: Coarse Perception, Cross Inst.: Cross-Instance Fine-grained Perception, Single-Inst.: Single-Instance Fine-grained Perception.
Category Model Logical Relation Attribute Coarse Cross-Inst. Single-Inst. Average
Base Qwen2.5-VL-3B 61.77 41.54 76.62 73.55 64.32 82.06 72.99
SSL4RL Rotation 65.84 80.54 83.89 80.21 71.53 84.76 80.38
Jigsaw 64.28 78.68 82.76 79.72 66.24 84.85 79.11
Contrastive 61.12 73.42 71.81 65.38 58.39 78.50 69.27
Position 67.65 77.19 82.22 82.15 66.51 85.39 80.08
Maximal Improvement $\uparrow$ 5.88 $\uparrow$ 39.00 $\uparrow$ 6.77 $\uparrow$ 8.60 $\uparrow$ 7.21 $\uparrow$ 3.33 $\uparrow$ 7.39
Table 2: Test performance (%) on SEED-Bench downstream tasks. TU: Text Understanding, VR: Visual Reasoning, SU: Scene Understanding, IId: Instance Identity, IIn: Instance Interaction, IA: Instance Attributes, IL: Instance Location, SR: Spatial Relation, IC: Instances Counting.
Category Model TU VR SU IId IIn IA IL SR IC Average
Base Qwen2.5-VL-3B 41.67 53.78 60.35 63.24 64.95 62.87 58.79 51.60 60.52 60.83
SSL4RL Rotation 45.24 73.41 73.65 72.80 67.01 71.03 61.76 54.03 64.12 69.10
Jigsaw 62.73 69.50 72.67 73.07 70.10 74.32 60.74 55.71 46.43 68.24
Contrastive 28.57 67.07 67.10 68.38 64.95 61.22 63.70 51.75 54.03 61.90
Position 52.38 70.69 73.56 72.75 62.89 72.51 64.62 55.25 64.20 69.77
Maximal Improvement $\uparrow$ 21.06 $\uparrow$ 19.63 $\uparrow$ 13.30 $\uparrow$ 9.83 $\uparrow$ 5.15 $\uparrow$ 9.64 $\uparrow$ 5.83 $\uparrow$ 4.11 $\uparrow$ 3.68 $\uparrow$ 8.94
Analysis. Our experimental analysis reveals that the Rotation and Position pretext tasks consistently yield the strongest performance gains. The success of the Position task is intuitive, as localizing a patch within the global image compels the model to integrate fine-grained local details with the overall scene layout, fostering integrated spatial understanding. In contrast, the remarkable effectiveness of the Rotation task presents a more nuanced insight. Despite its perceptual simplicity for humans, we find the task poses a considerable challenge to VLMs, as evidenced by the base model’s near-chance accuracy. We interpret this as evidence that rotation prediction facilitates “anti-commonsense” learning. The model’s pre-training instills a strong prior for canonical orientations, e.g., people are typically depicted upright. Rotated images violate this prior, forcing the model to reconcile anomalous inputs with its existing knowledge, thereby sharpening its relational reasoning and visual comprehension.
Conversely, the standard Contrastive task leads to performance degradation on some downstream tasks. We hypothesize that its default augmentations are insufficiently challenging, causing the model to overfit to superficial invariances without learning semantically meaningful structures. Subsequent ablation studies confirm this: employing a more aggressive augmentation strategy recovers significant performance. This finding highlights the critical need to align SSL task difficulty with model capacity to elicit generalizable feature learning, as explored in Section 4.2.2.
Observations. Through a qualitative analysis of model responses, we observe two key improvements attributable to SSL4RL training. (1) Sharper Attention: The trained models exhibit more precise attention alignment with text queries. For instance, when queried about “hair” (Figure 3, Left), the base model’s attention is diffuse, whereas our model accurately focuses on the relevant region. This is not merely an effect of sparsity, as evidenced by the “sky” query (Figure 3, Right): our model attends to the entire sky area, while the base model activates only sparse, scattered pixels. (2) Reduced Language Bias: SSL4RL mitigates over-reliance on linguistic priors, fostering greater dependence on visual evidence. For example, when asked about a chandelier’s color (Appendix Figure 8), the base model defaults to a common-sense response (e.g., a typical decorative color), while our model first localizes the object and then answers based on the actual appearance.
Figure 3: Cross-Attention Heatmap Comparison. More instances are shown in Appendix E.
η Takeaway SSL as Intrinsic Reward Sharpens VLM Reasoning. The SSL4RL paradigm demonstrably enhances vision-language reasoning by repurposing SSL tasks as intrinsic rewards. It deepens the perception and understanding of the image itself, leading towards more precise visual attention and less language bias. Task Choice is Critical. SSL tasks show effectiveness when their inherent semantic aligns with core reasoning skills (e.g, Position and Rotation), while an inappropriate task may induce negative transfer and hinder downstream performance.
4.2 Ablation Study
4.2.1 Task Difficulty Scaling
To investigate the role of task difficulty, we design more challenging self-supervised learning (SSL) tasks by intensifying their corruption strategies. Specifically, for the Position and Jigsaw tasks, we increase the crop granularity from 2 to 3, resulting in a finer 3×3 grid of patches. For the Contrastive task, we enhance the augmentation strength by raising the application probability from 0.2 to 0.8 and reducing the maximum crop scale from 1.0 to 0.3, thereby generating positive samples that differ more substantially from the anchor image. For the Rotation task, we refine the rotation angle interval from 90° to 45°, increasing the complexity of the angle prediction.
Figure 4: Test accuracy (%) on MMBench, varying the SSL task difficulty.
The impact of task difficulty varies considerably across different SSL tasks, as evidenced by the results in Figure 4 (full results in Appendix C). The performances for the Contrastive task shows a marked improvement upon increasing its difficulty, elevated from 69.27% to 77.89% on MMBench and from 61.90% to 65.00% on SEED-Bench. Conversely, the benefits are marginal for the Position task and even detrimental for Rotation and Jigsaw. A plausible explanation is that the Contrastive task’s inherent simplicity as a binary discrimination problem means that raising the difficulty compels the model to learn more robust and informative features. On the other hand, exacerbating the difficulty of already challenging tasks like Jigsaw might induce over-specialization to the SSL objective, resulting in a negative transfer where the learned representations are less transferable or even counterproductive for downstream vision-language reasoning.
4.2.2 Model Size Scaling
To investigate the scalability of the SSL4RL paradigm, we apply it to the larger Qwen2.5-VL-7B-Instruct model. The results (Table 3 and Appendix Table 8) confirm that SSL4RL still yields improvements over the base model, notably enhancing Logical Reasoning on MMBench by 2.57% and Text Understanding on SEED-Bench by 3.57%. However, the gains are less pronounced than those observed with the 3B model. Even after increasing the task difficulty to a 5×5 grid for Jigsaw and Position tasks, no significant improvement was observed (see Appendix C).
We attribute these diminishing returns to a fundamental ceiling effect imposed by the predefined SSL tasks. The absolute difficulty of the four SSL tasks is fixed, presenting a suitable challenge for the 3B model but potentially failing to engage the full capabilities of a 7B model. The larger model’s enhanced abilities could render the tasks trivial, thereby weakening their effectiveness as a learning signal. This underscores a key insight: the efficacy of an SSL task is contingent on its ability to present a non-trivial challenge commensurate with the model’s capacity. Consequently, a primary challenge for future work is the design of adaptive or inherently more complex SSL objectives that can continue to provide a learning signal for large-scale models.
Table 3: Test performance (%) of 7B models on MMBench downstream tasks.
Category Model Logical Relation Attribute Coarse Cross-Inst. Single-Inst. Average
Base Qwen2.5-VL-7B 76.49 84.68 85.69 84.66 84.49 89.15 86.37
SSL4RL-7B Rotation 78.70 86.16 87.13 85.91 88.47 88.92 87.50
Jigsaw 74.87 83.26 85.18 85.24 86.28 89.09 86.27
Contrastive 77.73 83.19 84.58 85.89 85.29 87.74 86.25
Position 79.06 82.13 85.29 85.02 85.42 88.95 86.25
Maximal Improvement $\uparrow$ 2.57 $\uparrow$ 1.48 $\uparrow$ 1.44 $\uparrow$ 1.25 $\uparrow$ 3.98 $\downarrow$ 0.06 $\uparrow$ 1.13
4.2.3 Task Combination
Figure 5: Test accuracy of SSL tasks.
The preceding sections primarily investigate the effect of individual SSL rewards. A natural subsequent question is whether combining them during training can yield better performance compared to any single reward. To explore this, we train the Qwen2.5-VL-3B-Instruct model using a combination of all four SSL rewards. As illustrated in Figure 5, the combined approach, somewhat counterintuitively, does not yield significant improvements over the best single-reward setups. We hypothesize that this lack of additive improvement stems from several potential factors. First, different SSL tasks may encourage the model to learn complementary yet potentially conflicting feature representations. Simultaneously optimizing for multiple, distinct objectives could create a difficult optimization landscape where the model struggles to find a unified representation that satisfies all constraints at once, leading to interference rather than synergy. Second, the individual reward signals might vary in scale and dynamics, making it non-trivial to balance their contributions effectively without careful reward shaping or weighting. A simple averaging of rewards might drown out the most useful learning signals. This finding suggests that a naive combination of SSL rewards is insufficient for achieving cumulative gains. It points to the need for more sophisticated integration strategies, such as dynamic reward weighting, curriculum learning that schedules different tasks, or even a meta-learner that selects the most beneficial task at different training stages.
η Takeaway Goldilocks Principle of Task Difficulty. The effectiveness of an SSL task is contingent on its difficulty being appropriately matched to the model’s capacity. Insufficient challenge provides a weak learning signal, while excessive difficulty leads to negative transfer. Diminishing Returns with Model Size. The performance gains from the four SSL tasks diminish as model size increases (from 3B to 7B), suggesting designing SSL tasks with inherently higher complexity for large-scale models. Non-additivity of Rewards. A naive combination of multiple SSL rewards does not yield cumulative improvements, indicating potential optimization conflicts and underscoring the need for sophisticated integration strategies rather than simple averaging.
4.3 Experiments on Vision-Centric Tasks
Benchmarks. We further evaluate the SSL4RL paradigm on vision-centric tasks using the ImageNet-1K dataset (Deng et al., 2009), which comprises approximately 1.3 million images across 1,000 categories. From this dataset, we construct a balanced subset of 100,000 training and 10,000 test images. To probe reasoning capabilities at varying difficulty levels, we design three question types: (1) Completion: directly answer with the species name; (2) Choice-20: select the correct species from 20 candidates; (3) Choice-200: select the correct species from 200 candidates.
Results. As shown in Table 4, models fine-tuned with SSL4RL consistently outperform the base model across all question types on the ImageNet-1K classification task. Consistent with findings on reasoning benchmarks, the Position task leads to the largest performance gains, e.g., 67.14% vs. 57.20% on Choice-200. However, a key divergence emerges with the Contrastive task. While it underperformed on vision-language reasoning, it shows competitive results on ImageNet classification. We attribute this result to the nature of the downstream benchmark. As an instance discrimination task, ImageNet benefits from learning strong semantic representations through invariance to augmentations—precisely the strength of contrastive learning. This result also verifies that task selection for SSL4RL must consider the specific capabilities required by the target application.
Table 4: Test performance (%) on ImageNet downstream tasks.
Category Model Completion Choice-20 Choice-200
Base Model Qwen2.5-VL-3B 24.93 85.22 57.10
SSL4RL Rotation 29.19 87.26 58.48
Jigsaw 28.08 87.07 60.80
Contrastive 26.84 89.51 61.78
Position 28.76 92.35 67.14
5 Extension to Other Domains: An Example on Graph
Having established SSL4RL for vision-language reasoning, we now explore its broader applicability. The paradigm’s core principle—generating rewards from data transformations—naturally extends to any domain with rich structural information. Beyond images, graph-structured data presents a compelling candidate, given its explicit relational semantics that are amenable to various pretext tasks. In this section, we empirically validate this potential by adapting SSL4RL to the graph domain.
Tasks and Reward Definitions. We introduce three graph-based SSL tasks, defined as follows: (1) Attribute Mask (Jin et al., 2020; Hu et al., 2019): A subset of node descriptions is randomly masked. The reward is quantified by the model’s accuracy in reconstructing the original masked features. (2) Neighbor Prediction (Kipf & Welling, 2016): For a target node within a partially observed graph, the model is rewarded for correctly identifying its adjacent nodes. (3) Link Prediction (Hu et al., 2020; Hou et al., 2022): Given a pair of nodes and a partial graph structure, the model receives a reward for accurately classifying the presence or absence of an edge connecting them.
Benchmarks. We evaluate our method on benchmark datasets curated from TAGLAS (Feng et al., 2024), a comprehensive collection of text-attributed graphs. The evaluation encompasses two key tasks: (1) Node-level classification on the Cora and PubMed co-citation graphs and the WikiCS page relation graph; (2) Link-level prediction on the Products co-purchase graph and the fb15k237 and wn18rr knowledge graphs.
Results and Observations. The results in Table 5 demonstrate the successful application of SSL4RL to graph-structured data. The 3B model shows substantial improvements, with gains up to 13.79% on average, while the 7B model exhibits diminishing returns, mirroring our observations in the visual domain and reinforcing the “difficulty-capacity matching” principle. Furthermore, the relative effectiveness of the self-supervised tasks is contingent upon the nature of the downstream objective. Tasks emphasizing structural reasoning (Link and Neighbor Prediction) yield better performance on relation-centric tasks such as link prediction. Conversely, tasks focused on feature reconstruction (Attribute Mask) demonstrate a comparative advantage on node classification benchmarks. These findings not only validate the generalizability of the SSL4RL framework beyond the visual modality but also highlight the critical importance of aligning the pretext task’s inductive bias with the target application.
Table 5: Test performance (%) on downstream graph tasks.
Category Model Cora PubMed WikiCS Products fb15k237 wn18rr Average
Base model Qwen2.5-VL-3B 21.80 64.26 30.50 7.93 26.30 29.80 30.09
SSL4RL-3B Attribute 55.80 73.27 57.62 8.03 32.50 36.10 43.88
Neighbor 39.00 74.37 50.84 12.65 36.10 36.60 41.59
Link 31.30 71.97 55.93 4.91 46.50 41.10 41.95
Maximal Improvement $\uparrow$ 34.00 $\uparrow$ 10.11 $\uparrow$ 27.12 $\uparrow$ 4.42 $\uparrow$ 20.20 $\uparrow$ 11.30 $\uparrow$ 13.79
Base model Qwen2.5-7B-Instruct 64.80 69.86 49.15 50.50 30.10 32.50 57.73
SSL4RL-7B Attribute 63.80 59.55 54.23 51.30 37.00 32.70 49.76
Neighbor 63.10 70.87 52.54 43.47 33.00 34.00 49.49
Link 67.70 72.27 55.93 48.89 21.30 39.50 50.93
Maximal Improvement $\uparrow$ 2.90 $\uparrow$ 2.41 $\uparrow$ 6.78 $\uparrow$ 0.80 $\uparrow$ 6.90 $\uparrow$ 7.00 $\uparrow$ 1.45
6 Conclusions
We have introduced SSL4RL, a framework that repurposes self-supervised tasks as verifiable reinforcement learning rewards for post-training vision–language models. Our study shows that SSL4RL not only improves performance on vision-centric benchmarks such as ImageNet-1K, but also enhances multimodal reasoning, achieving substantial gains on MMBench and SEED-Bench. These findings suggest a broader principle: verifiable and scalable supervision signals are already embedded in self-supervision, and with proper task selection they can drive alignment and reasoning in VLMs without reliance on external verifiers, judges, or costly human labels. Looking forward, SSL4RL opens a path toward safer and more capable multimodal foundation models by unifying the strengths of self-supervision and reinforcement learning.
Ethics Statement
We are not aware of any specific ethical concerns related to this work. All experiments are conducted on publicly available or synthetic datasets, without the use of sensitive or proprietary information.
Reproducibility Statement
We provide complete details of our methods, hyperparameters, datasets, and evaluation metrics in both the main paper and the appendix. To further support transparency and reproducibility, we will release our code upon acceptance.
Appendix A Detailed Results of MMBench
In this section, we present the detailed leaf task results of MMBench in Tables 6 and 7.
Table 6: Test performance (%) of 3B models on MMBench downstream tasks. IR: Identity Reasoning, PPR: Physical Property Reasoning, FR: Function Reasoning, OL: Object Localization, SIU: Structuralized Imagetext Understanding, AtR: Attribute Recognition, FP: Future Prediction, SR: Spatial Relationship, IS: Image Scene, IQ: Image Quality, Ace: Action Recognition, AC: Attribute Comparison, IT: Image Topic, NR: Nature Relation, PR: Physical Relation, SR: Social Relation, CR: Celebrity Recognition, IS: Image Style, OCR: OCR, IE: Image Emotion.
Category Model IR PPR FR OL SIU AtR FP SR IS IQ Ace AC IT NR PR SR CR IS OCR IE Average
Base Qwen2.5-VL-3B 86.36 60.27 83.22 63.17 71.99 81.44 51.54 45.20 91.15 46.67 84.65 63.12 82.86 64.25 47.87 38.37 89.39 82.55 94.23 64.50 72.99
SSL4RL-3B Rotation 97.16 66.67 87.83 63.49 75.53 85.23 56.15 56.50 95.82 55.33 89.30 68.79 87.14 70.39 64.89 88.37 95.45 87.26 94.87 75.50 80.38
Jigsaw 96.59 64.84 86.84 62.86 71.63 88.26 56.92 42.94 97.30 52.67 86.98 68.79 85.00 67.60 62.77 86.63 94.70 89.15 93.59 74.50 79.11
Contrastive 90.34 52.05 73.03 60.32 69.15 78.03 53.08 32.77 88.45 42.00 81.40 60.99 82.14 36.87 57.45 81.40 90.40 61.79 85.26 52.50 69.27
Position 95.45 64.38 86.84 66.03 78.37 86.36 56.92 42.37 96.31 58.00 89.77 67.38 86.43 68.16 60.64 85.47 94.95 91.51 94.23 78.50 80.08
Combination 95.45 61.64 82.24 68.25 77.66 89.77 54.62 45.76 94.10 55.33 90.23 63.83 87.86 61.45 54.26 90.12 94.19 89.15 94.87 67.50 78.77
Table 7: Test performance (%) of 7B models on MMBench downstream tasks. IR: Identity Reasoning, PPR: Physical Property Reasoning, FR: Function Reasoning, OL: Object Localization, SIU: Structuralized Imagetext Understanding, AtR: Attribute Recognition, FP: Future Prediction, SR: Spatial Relationship, IS: Image Scene, IQ: Image Quality, Ace: Action Recognition, AC: Attribute Comparison, IT: Image Topic, NR: Nature Relation, PR: Physical Relation, SR: Social Relation, CR: Celebrity Recognition, IS: Image Style, OCR: OCR, IE: Image Emotion.
Category Model IR PPR FR OL SIU AtR FP SR IS IQ Ace AC IT NR PR SR CR IS OCR IE Average
Base Qwen2.5-VL-7B 98.30 66.67 92.11 74.60 82.98 88.64 70.00 76.27 97.05 58.00 92.09 85.11 91.43 88.83 69.15 92.44 97.22 94.81 96.15 82.00 86.37
SSL4RL-7B Rotation 97.73 71.23 92.43 71.11 89.72 90.15 67.69 78.53 97.79 61.33 92.56 94.33 90.71 85.47 71.28 93.60 96.97 96.23 97.44 83.50 87.50
Jigsaw 97.73 64.38 93.42 75.24 80.50 88.26 69.23 77.40 97.30 61.33 93.49 87.94 90.71 87.15 64.89 92.44 96.72 93.87 96.15 83.00 86.27
Contrastive 98.30 63.01 92.43 74.29 85.46 85.61 70.00 77.97 97.30 63.33 93.49 84.40 88.57 85.47 67.02 91.28 97.47 95.75 93.59 84.50 86.25
Position 98.30 63.01 92.43 74.29 85.46 85.61 70.00 77.97 97.30 63.33 93.49 84.40 88.57 85.47 67.02 91.28 97.47 95.75 93.59 84.50 86.25
Combination 97.73 73.97 92.11 71.43 90.43 95.08 70.77 74.58 97.79 61.33 93.49 81.56 92.14 89.39 68.09 93.02 97.47 93.87 96.79 84.50 85.78
Appendix B Results of SSL4RL 7B-model on SEED-Bench
In Table 8, we present the SEED-Bench results for the SSL4RL 7B-models.
Table 8: Test performance (%) of 7B models on SEED-Bench downstream tasks. TU: Text Understanding, VR: Visual Reasoning, SU: Scene Understanding, IId: Instance Identity, IIn: Instance Interaction, IA: Instance Attributes, IL: Instance Location, SR: Spatial Relation, IC: Instances Counting.
Category Model TU VR SU IId IIn IA IL SR IC Average
Base Qwen2.5-VL-7B 72.62 77.95 77.99 77.44 75.26 76.19 71.98 62.56 69.55 74.70
SSL4RL-7B Rotation 76.19 78.25 78.59 77.94 73.20 76.90 73.11 64.23 69.76 75.33
Jigsaw 68.49 76.64 77.58 78.21 73.20 79.76 71.78 62.10 67.86 74.64
Contrastive 71.43 78.85 78.28 78.32 76.29 76.47 72.70 64.69 70.33 75.27
Position 70.24 80.66 78.82 77.44 71.13 77.97 72.19 64.38 69.39 75.56
Maximal Improvement $\uparrow$ 3.57 $\uparrow$ 2.71 $\uparrow$ 0.83 $\uparrow$ 0.88 $\uparrow$ 1.03 $\uparrow$ 1.78 $\uparrow$ 3.57 $\uparrow$ 2.13 $\uparrow$ 0.78 $\uparrow$ 0.86
Appendix C Detailed Results of the Ablation Study about Difficulty
From Table 9 to Table 12, we provide detailed experimental results for the ablation study on task difficulty of MMBench and SEED-Bench.
Table 9: Test performance (%) of 3B models trained with different task difficulties on MMBench downstream tasks. IR: Identity Reasoning, PPR: Physical Property Reasoning, FR: Function Reasoning, OL: Object Localization, SIU: Structuralized Imagetext Understanding, AtR: Attribute Recognition, FP: Future Prediction, SR: Spatial Relationship, IS: Image Scene, IQ: Image Quality, Ace: Action Recognition, AC: Attribute Comparison, IT: Image Topic, NR: Nature Relation, PR: Physical Relation, SR: Social Relation, CR: Celebrity Recognition, IS: Image Style, OCR: OCR, IE: Image Emotion.
Model Difficulty IR PPR FR OL SIU AtR FP SR IS IQ Ace AC IT NR PR SR CR IS OCR IE Average
Qwen2.5-VL-3B - 86.36 60.27 83.22 63.17 71.99 81.44 51.54 45.20 91.15 46.67 84.65 63.12 82.86 64.25 47.87 38.37 89.39 82.55 94.23 64.50 72.99
Rotation 90-degree 97.16 66.67 87.83 63.49 75.53 85.23 56.15 56.50 95.82 55.33 89.30 68.79 87.14 70.39 64.89 88.37 95.45 87.26 94.87 75.50 80.38
45-degree 93.75 64.84 87.83 66.67 74.11 90.15 58.46 46.33 95.33 53.33 88.84 65.25 87.14 65.92 54.26 90.12 93.94 86.79 92.31 70.50 79.16
Jigsaw 2x2 96.59 64.84 86.84 62.86 71.63 88.26 56.92 42.94 97.30 52.67 86.98 68.79 85.00 67.60 62.77 86.63 94.70 89.15 93.59 74.50 79.11
3x3 97.73 63.93 88.82 62.54 75.18 87.50 55.38 59.89 96.07 50.67 86.51 74.47 86.43 74.86 50.00 89.53 94.19 88.21 96.15 78.00 80.38
Contrastive Weak 90.34 52.05 73.03 60.32 69.15 78.03 53.08 32.77 88.45 42.00 81.40 60.99 82.14 36.87 57.45 81.40 90.40 61.79 85.26 52.50 69.27
Strong 94.89 64.84 82.24 63.49 77.30 86.36 51.54 44.07 95.82 47.33 89.30 70.21 85.71 59.78 52.13 86.63 94.44 87.74 92.31 70.50 77.89
Position 2x2 95.45 64.38 86.84 66.03 78.37 86.36 56.92 42.37 96.31 58.00 89.77 67.38 86.43 68.16 60.64 85.47 94.95 91.51 94.23 78.50 80.08
3x3 97.16 63.47 87.50 66.35 75.53 88.64 60.00 52.54 95.33 58.67 86.51 76.60 88.57 70.39 58.51 88.95 95.96 92.45 93.59 77.50 81.03
Table 10: Test performance (%) of 3B models trained with different task difficulties on SEED-Bench downstream tasks. IC: Instances Counting, IA: Instance Attributes, SU: Scene Understanding, IId: Instance Identity, IIn: Instance Interaction, VR: Visual Reasoning, IL: Instance Location, SR: Spatial Relation, TU: Text Understanding.
Model Difficulty IC IA SU IId IIn VR IL SR TU Average
Qwen2.5-VL-3B - 60.52 62.87 60.35 63.24 64.95 53.78 58.79 51.60 41.67 60.83
Rotation 90-degree 64.12 71.03 73.65 72.80 67.01 73.41 61.76 54.03 45.24 69.10
45-degree 60.60 69.26 72.70 72.53 68.04 72.81 63.60 55.25 38.10 67.81
Jigsaw 2x2 62.73 69.50 72.67 73.07 70.10 74.32 60.74 55.71 46.43 68.24
3x3 62.93 68.94 73.27 72.69 69.07 74.32 62.58 52.97 35.71 68.10
Contrastive Weak 54.03 61.22 67.10 68.38 64.95 67.07 63.70 51.75 28.57 61.90
Strong 57.54 64.75 70.68 70.56 67.01 70.39 63.39 55.86 29.76 65.00
Position 2x2 64.20 72.51 73.56 72.75 62.89 70.69 64.62 55.25 52.38 69.77
3x3 61.87 72.53 73.84 73.89 69.07 74.02 64.62 58.30 44.05 69.80
Table 11: Test performance (%) of 7B models trained with different task difficulties on MMBench downstream tasks. IR: Identity Reasoning, PPR: Physical Property Reasoning, FR: Function Reasoning, OL: Object Localization, SIU: Structuralized Imagetext Understanding, AtR: Attribute Recognition, FP: Future Prediction, SpR: Spatial Relationship, ISc: Image Scene, IQ: Image Quality, Ace: Action Recognition, AC: Attribute Comparison, IT: Image Topic, NR: Nature Relation, PR: Physical Relation, SoR: Social Relation, CR: Celebrity Recognition, ISt: Image Style, OCR: OCR, IE: Image Emotion.
Model Difficulty IR PPR FR OL SIU AtR FP SpR ISc IQ Ace AC IT NR PR SoR CR ISt OCR IE Average
Base Qwen2.5-VL-7B 98.30 66.67 92.11 74.60 82.98 88.64 70.00 76.27 97.05 58.00 92.09 85.11 91.43 88.83 69.15 92.44 97.22 94.81 96.15 82.00 86.37
Jigsaw 2x2 97.73 64.38 93.42 75.24 80.50 88.26 69.23 77.40 97.30 61.33 93.49 87.94 90.71 87.15 64.89 92.44 96.72 93.87 96.15 83.00 86.27
3x3 97.16 68.04 92.76 72.38 82.27 90.53 72.31 75.14 97.54 60.67 92.09 86.52 88.57 92.18 69.15 91.28 96.97 95.28 98.08 84.50 86.74
4x4 97.73 72.60 95.07 75.24 87.59 92.80 70.00 76.27 97.30 58.00 92.56 87.23 92.14 88.83 67.02 91.28 97.47 94.34 98.08 81.00 87.52
Position 2x2 98.30 63.01 92.43 74.29 85.46 85.61 70.00 77.97 97.30 63.33 93.49 84.40 88.57 85.47 67.02 91.28 97.47 95.75 93.59 84.50 86.25
3x3 95.45 68.49 93.09 77.46 88.30 94.70 73.08 77.40 97.79 62.67 94.42 89.36 90.71 87.15 67.02 89.53 97.47 94.34 97.44 81.50 85.87
4x4 97.73 64.84 94.74 74.92 88.65 95.45 73.08 72.32 97.79 58.00 94.42 82.98 89.29 88.83 69.15 90.12 98.23 94.34 98.08 82.50 85.27
Table 12: Test performance (%) of 7B models trained with different task difficulties on SEED-Bench downstream tasks. IC: Instances Counting, IA: Instance Attributes, SU: Scene Understanding, IId: Instance Identity, IIn: Instance Interaction, VR: Visual Reasoning, IL: Instance Location, SR: Spatial Relation, TU: Text Understanding.
Category Model TU VR SU IId IIn IA IL SR IC Average
Base Qwen2.5-VL-7B 72.62 77.95 77.99 77.44 75.26 76.19 71.98 62.56 69.55 74.70
Jigsaw 2x2 68.49 76.64 77.58 78.21 73.20 79.76 71.78 62.10 67.86 74.64
3x3 70.41 77.26 78.50 78.32 76.29 79.15 72.60 63.62 70.24 75.54
4x4 70.04 76.70 77.23 77.39 74.23 79.15 72.29 63.62 70.24 74.85
Position 2x2 70.24 80.66 78.82 77.44 71.13 77.97 72.19 64.38 69.39 75.56
3x3 68.82 77.93 78.53 78.75 74.23 78.55 73.01 64.08 76.19 74.45
4x4 68.98 77.41 77.58 77.33 76.29 78.55 71.47 61.64 76.19 73.94
Appendix D RL Training Reward Curves
Figure 6 illustrates the evolving reward signals of different SSL tasks during reinforcement learning training. Despite being intuitive for humans, these SSL tasks pose significant challenges for the base VLM. For instance, Qwen2.5-VL-3B initially achieves only about 20% accuracy on the Rotation task and nearly zero on the Jigsaw task, highlighting its limited low-level perceptual capability regarding spatial and structural image properties. Position and Rotation tasks exhibit gradual learning curves, plateauing at a reward of about 0.8 after approximately 200 epochs. For the harder Jigsaw task, the reward surges and plateaus within 20 epochs for 3B models, while 7B models present more smooth reward curves, indicating a stronger learning capability of large-scale models. For the Contrastive task, models rapidly surge to 1.0 within 50 epochs due to the task’s easy complexity. Figure 7 presents the reward curves of tasks with different difficulties. For the same task, increasing the difficulty slows down the learning process and lowers the final converging rewards, showing the sensitivity of the models to tasks’ difficulty.
(a) MMBench (3B)
(b) SEED-Bench (3B)
(c) MMBench (7B)
(d) SEED-Bench (7B)
Figure 6: Reward curves of SSR4RL models during reinforcement learning.
(a) Rotation
(b) Jigsaw
(c) Contrastive
(d) Position
Figure 7: Reward curves of SSL4RL 3B-models on MMBench, comparing different difficulties.
Appendix E Case Analysis and Visualization
In Figure 8, we show that SSL4RL mitigates over-reliance on linguistic priors, fostering greater dependence on visual evidence. For example, when asked about a chandelier’s color, the base model defaults to a common-sense response (e.g., a typical decorative color), while our model first localizes the object and then answers based on the actual appearance.
Figure 8: Instances of VLMs’ loss on image information. After receiving textual instructions, VLMs may be more inclined to rely on the encoded textual knowledge for reasoning rather than carefully observing the content of the image.
Besides, we visualize the attention maps of the baseline model, i.e., Qwen2.5-VL-3B and our models on several examples from the SEED-Bench dataset(Li et al., 2023). We pick a dominant token from the questions of each example, calculate the attention map of the first generated token to that input token, and average the attention matrices of all heads and all layers of the language model. The results in Figure 9 illustrate that our models consistently display more focused attention towards the regions in the images corresponding to the selected token, which indirectly proves the superior performance of our models.
Figure 9: Comparisons of Attention Maps.
Appendix F SSL Task Examples
In this section, we show a specific instance of Rotation, Jigsaw, Contrastive, and Position tasks to illustrate the SSL task design.
Rotation Example Query: These are two images. The second image is a rotated version of the first image. Please determine how many degrees the second image has been rotated counter-clockwise relative to the first image. You must reason step-by-step and then provide the final answer. The output must strictly follow this format: <think>your reasoning here </think><answer>number_of_degrees</answer>. Answer: 270 [Uncaptioned image] [Uncaptioned image]
Jigsaw Example Query: <image><image><image><image> <image><image><image><image><image> The provided images represent 9 parts of an original image, divided into a 3x3 grid. Your task is to determine the correct order of these parts to reconstruct the original image. Starting from the top-left corner, proceed row by row, from left to right and top to bottom, to arrange the parts. The output should be a string of numbers, separated by a comma, where each number corresponds to the original position of the patches in the restored image. For instance, “3,1,9,2,8,5,4,6,7” would indicate the positions of the patches in the correct order. Before providing the final result, you must reason through the puzzle step by step. Consider the relative placement of each part and how they fit together. Your answer should strictly follow this format: <think>your step-by-step reasoning here</think><answer>order</answer> Answer: 2,7,6,1,3,5,9,8,4 [Uncaptioned image]
Contrastive Example Query: <image><image> The provided images are augmentations of the same original image or two different images. The augmentations may include random cropping, color adjustments, grayscale conversion, blurring, and flipping. Please think step-by-step and determine if these two images are possibly derived from the same original image. If the provided images are from the same original image, respond with “positive”; if they correspond to different original images, respond with “negative”. Your answer should strictly follow this format: <think>your step-by-step reasoning here</think><answer>positive/negative</answer> Answer: positive [Uncaptioned image] [Uncaptioned image]
Position Example Query: <image><image> The second image in an augmented version of a crop in the first image. The augmentations may include grayscale, color jitter, solarization, etc. Please determine which part of the first image the second image is from. The second image is partitioned into 3x3 parts, and the first image can be only from one of the parts, but cannot be across two parts. The answer should be in the format of x/y, where x is the row number (from top to bottom) and y is the column number (from left to right). For example, 1/1 indicates the top-left part, and 1/3 indicates the top-right part. Both x and y may take values from 1 to 3. Your answer should strictly follow this format: <think>your step-by-step reasoning here</think><answer>x/y</answer> Answer: 3/3 [Uncaptioned image] [Uncaptioned image]
Appendix G Downstream Benchmark Example
In this section, we show a specific instance of Rotation, Jigsaw, Contrastive, and Position tasks to illustrate the SSL task design.
Imagenet-Completion Example Query: <image>This is an image containing an object. Please identify the species of the object based on the image. The output answer format should be as follows: <think>… </think><answer>species name</answer>. Please strictly follow the format. Answer: tench, Tinca, tinca [Uncaptioned image]
Imagenet-Choice10 Example Query: <image>This is an image containing an object. Please identify the species of the object based on the image. The output answer format should be as follows: <think>… </think><answer>species name</answer> Please strictly follow the format. Please select the correct species name from the following options: Ursus americanus, shoe shop, brush wolf, essence, malemute, scoreboard, tench, ruddy turnstone, Salamandra salamandra, koala. Answer: tench [Uncaptioned image]
MMBench Example Query: <image>Identify the question that Madelyn and Tucker’s experiment can best answer. A. Does Madelyn’s snowboard slide down a hill in less time when it has a thin layer of wax or a thick layer of wax? B. Does Madelyn’s snowboard slide down a hill in less time when it has a layer of wax or when it does not have a layer of wax? C. NaN. D. NaN. Answer: B [Uncaptioned image]
SEED-Bench Example Query: <image>How many towels are in the image? A. One. B. Two. C. Three. D. Four. Answer: A [Uncaptioned image]